# 1]Ling Team, Ant Group

### v1

@@ 1 Introduction
Figure 1: Comparison between WSM and mainstream learning rate scheduling strategies.
@@ 3.1 Theoretical Connection Between LR Decay and Checkpoint Merging
The core idea of checkpoint merging in this work is to take an ordered list of checkpoints, $[\theta_{n},\theta_{n+1},\dots,\theta_{n+k}]$, and apply a merge function to generate a single model $\hat{\theta}_{n+k}$. Here, $\theta_{i}\in\mathbb{R}^{d}$ represents the model’s parameter vector at the $i$-th training iteration. The most general form is a weighted average of the checkpoints:
@@
Figure 2: Visualization of checkpoint merging weight distributions and their corresponding decay functions over a merge duration of 10 checkpoints (larger checkpoint index denotes more recent checkpoints). (a) Exponential Moving Average (EMA) weights, exhibiting convex decay characteristics; (b) Uniform averaging weights, demonstrating linear decay behavior; (c) and (d) Our theoretically derived weights, designed to approximate cosine and 1-sqrt decay patterns. Note: We draw the decay curves smoothly for better illustration, although they are actually step-like due to their discrete nature.
@@
This shows that a weight coefficient of $\sum_{j=i}^{k}c_{j}$ is applied to the gradient update $g_{n+i-1}$ from step $n+i-1$. By defining a new set of weights for the gradient updates, $w_{i}=\sum_{j=i}^{k}c_{j}$, we arrive at the final equivalent expression for checkpoint merging:
@@ 3.2 Emulating LR Decay through Checkpoint Merging
$lr(t)=\begin{cases}lr_{peak}\cdot\frac{t}{T_{warmup}}&\text{if }t<T_{warmup}\\ lr_{peak}&\text{if }t\geq T_{warmup}\\ \end{cases}$
@@ 4.3.2 Impact of Merging Algorithm
As derived in Section 3.1, the checkpoint merging process can be viewed as an approximation of a LR decay schedule, where the weighting scheme of the merge is analogous to the functional form of the decay curve. For instance, a simple mean average is analogous to a linear decay curve. An Exponential Moving Average (EMA) would correspond to a convex exponential decay curve. Our prior experiments with the WSD schedule revealed a performance hierarchy among decay curves: concave schedules (e.g., inverse square root) and linear schedules outperform convex schedules (details are provided in Appendix C). Building on these findings, we hypothesize that a merging algorithm designed to approximate decay curves that have been proven effective in WSD scheduling will similarly yield better results.
@@ 4.3.5 Impact on MoE Load Balancing
[…]
@@ 5.2 Model Merging
Model merging (Izmailov et al., 2018; Wortsman et al., 2022) has emerged as an efficient paradigm for model construction. This approach achieves effective knowledge transfer and performance improvement through parameter-level integration of multiple models. Model merging is primarily utilized in two distinct scenarios: (1) The integration of knowledge and capabilities from multiple independently trained models into a single parameter set, with the objective of preserving maximal performance from each specialized model (Aakanksha et al., 2025; Ramé et al., 2024); and (2) the merging of checkpoints along a single training trajectory, which functions as a form of Polyak averaging (Polyak and Juditsky, 1992). In this application, the merging operation acts as a smoothing mechanism that reduces noise inherent in stochastic gradient-based optimization (Sanyal et al., 2023; Liu et al., 2024a; Kaddour, 2022; Li et al., 2023; Sandler et al., 2023). While Li et al. (2025) show model merging can achieve performance competitive with decay-based schedule, these techniques have also demonstrated practical utility in industrial-scale LLM development (Grattafiori et al., 2024; DeepSeek-AI et al., 2024; Aakanksha et al., 2025). Our work falls into the second category of merging checkpoints along a single trajectory. We establish a theoretical connection between this operation and learning rate decay and provide a principled approach to convert various LR decay strategies into a theoretically approximate model averaging implementation.
@@ Appendix A Experimental Settings
$n_{layers}$ $d_{model}$ $d_{ffn}$ $d_{expert}$ $n_{heads}$ $n_{kv\_head}$ $E$ $E_{a}$ $E_{s}$ $N$ $N_{a}$
@@
We use the AdamW optimizer (Loshchilov and Hutter, 2019) with hyperparameters set as follows: $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay of 0.1. Gradient clipping (Zhang et al., 2020) norm is set to 1.0. According to the scaling laws for MoE optimal hyper-parameters, the maximum learning rates were set to $3.74e{-4}$. The batch size is set to 2048, and with a maximum sequence length of 8K, each training batch contains 16M tokens.
@@ Appendix B Details of Theoretical Connection Between Decay and Merging
The core idea of checkpoint merging in this work is to take an ordered list of checkpoints, $[\theta_{n},\theta_{n+1},\dots,\theta_{n+k}]$, and apply a merge function to generate a single model $\hat{\theta}_{n+k}$. Here, $\theta_{i}\in\mathbb{R}^{d}$ represents the model’s parameter vector at the $i$-th training iteration.
@@
This shows that a weight coefficient of $\sum_{j=i}^{k}c_{j}$ is applied to the gradient update $g_{n+i-1}$ from step $n+i-1$. By defining a new set of weights for the gradient updates, $w_{i}=\sum_{j=i}^{k}c_{j}$, we arrive at the final equivalent expression for checkpoint merging:

### v2

@@ Abstract
Figure 1: Comparison between WSM (Warmup-Stable and Merge) and mainstream learning rate scheduling strategies. By leveraging checkpoint merging, WSM eliminates the learning rate decay phase and maintains a constant learning rate after warmup, enabling a fully autonomous and continuous training process.
@@ 3.1 Theoretical Connection Between LR Decay and Checkpoint Merging
The core idea of checkpoint merging in this work is to take an ordered list of checkpoints, $\theta_{n},\theta_{n+1},\dots,\theta_{n+k}]$, and apply a merge function to generate a single model $\hat{\theta}_{n+k}$. Here, $\theta_{i}\in\mathbb{R}^{d}$ represents the model’s parameter vector at the $i$-th training iteration. The most general form is a weighted average of the checkpoints:
@@
Figure 2: Visualization of checkpoint merging weight distributions and their corresponding decay functions over a span of 10 checkpoints (larger checkpoint indices denote more recent checkpoints). (a) Exponential Moving Average (EMA) weights, exhibiting convex decay characteristics; (b) Uniform averaging weights, demonstrating linear decay behavior; (c) and (d) Our theoretically derived weights, designed to approximate cosine and 1-sqrt decay patterns. Note: Decay curves are rendered smoothly for clarity, though the underlying weights are applied discretely at each checkpoint.
@@
This shows that a weight coefficient of $\sum_{j=i}^{k}c_{j}$ is applied to the gradient update $g_{n+i-1}$ from step n+i1n+i-1. By defining a new set of weights for the gradient updates, $w_{i}=\sum_{j=i}^{k}c_{j}$, we arrive at the final equivalent expression for checkpoint merging:
@@ 3.2 Emulating LR Decay through Checkpoint Merging
[…]
@@ 4.3.2 Impact of Merging Algorithm
As derived in Section 3.1, the checkpoint merging process can be viewed as an approximation of a LR decay schedule, where the weighting scheme of the merge is analogous to the functional form of the decay curve. For instance, a simple mean average is analogous to a linear decay curve. An Exponential Moving Average (EMA) would correspond to a convex exponential decay curve. Existing works (Hägele et al., 2024) and our prior experiments with the WSD schedule revealed a performance hierarchy among decay curves: concave schedules (e.g., inverse square root) and linear schedules outperform convex schedules (details are provided in Appendix C). Building on these findings, we hypothesize that a merging algorithm designed to approximate decay curves that have been proven effective in WSD scheduling will similarly yield better results.
@@ 4.3.5 Impact on MoE Load Balancing
We provide a more in-depth analysis of the implications of WSM on MoE router balance in Table 5. Specifically, the violation for a single expert is calculated as its relative deviation from the average load within its layer [^1] For a single layer, let $L$ be the vector of token loads for its experts and $\mu=\text{mean}(L)$, $\text{max\_violation}=\frac{\lvert\max(L)-\mu\rvert}{\mu}$ and $\text{min\_violation}=\frac{\lvert\min(L)-\mu\rvert}{\mu}$. The mean_global_max_violation represents the average of the highest violations across all layers (measuring the severity of “overloaded” experts), while mean_global_min_violation averages the violations for the least-utilized experts (measuring the risk of “routing collapse”). When comparing the merged checkpoint with a decayed checkpoint, the merged checkpoint achieves more balanced routing, although its loss is slightly higher. We argue that this trade-off—a marginally higher loss for superior downstream performance—is indicative of enhanced generalization rather than overfitting to the training data. More analysis of the parameter distribution is provided in Appendix D.
@@ 5.2 Model Merging
Model merging (Izmailov et al., 2018; Wortsman et al., 2022) has emerged as an efficient paradigm for model construction. This approach achieves effective knowledge transfer and performance improvement through parameter-level integration of multiple models. Model merging is primarily utilized in two distinct scenarios: (1) The integration of knowledge and capabilities from multiple independently trained models into a single parameter set, with the objective of preserving maximal performance from each specialized model (Aakanksha et al., 2025; Ramé et al., 2024); and (2) the merging of checkpoints along a single training trajectory, which functions as a form of Polyak averaging (Polyak and Juditsky, 1992). In this application, the merging operation acts as a smoothing mechanism that reduces noise inherent in stochastic gradient-based optimization (Sanyal et al., 2023; Liu et al., 2024a; Kaddour, 2022; Li et al., 2023; Sandler et al., 2023; Hägele et al., 2024). While Li et al. (2025) show model merging can achieve performance competitive with decay-based schedule, these techniques have also demonstrated practical utility in industrial-scale LLM development (Grattafiori et al., 2024; DeepSeek-AI et al., 2024; Aakanksha et al., 2025). Our work falls into the second category of merging checkpoints along a single trajectory. We establish a theoretical connection between this operation and learning rate decay and provide a principled approach to convert various LR decay strategies into a theoretically approximate model averaging implementation.
@@ Appendix A Experimental Settings
$n_{layers}$ $d_{model}$ $d_{ffn}$ $d_{expert}$ $n_{heads}$ $n_{kv\_head}$ $E$ $E_{a}$ $E_{s}$ NN $N_{a}$
@@
We use the AdamW optimizer (Loshchilov and Hutter, 2019) with hyperparameters set as follows: $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay of 0.1. Gradient clipping (Zhang et al., 2020) norm is set to 1.0. According to the scaling laws for MoE optimal hyper-parameters, the maximum learning rates were set to 3.74e43.74e{-4}. The batch size is set to 2048, and with a maximum sequence length of 8K, each training batch contains 16M tokens.
@@ Appendix B Details of Theoretical Connection Between Decay and Merging
The core idea of checkpoint merging in this work is to take an ordered list of checkpoints, $\theta_{n},\theta_{n+1},\dots,\theta_{n+k}]$, and apply a merge function to generate a single model $\hat{\theta}_{n+k}$. Here, $\theta_{i}\in\mathbb{R}^{d}$ represents the model’s parameter vector at the $i$-th training iteration.
@@
This shows that a weight coefficient of $\sum_{j=i}^{k}c_{j}$ is applied to the gradient update $g_{n+i-1}$ from step n+i1n+i-1. By defining a new set of weights for the gradient updates, $w_{i}=\sum_{j=i}^{k}c_{j}$, we arrive at the final equivalent expression for checkpoint merging: